A square has an area of 81 square inches. find the length of the sides, and the length of the diagonal
Answer by TimothyLamb(4379)

side length:
---
aa = 81
a = sqrt( 81 )
a = 9 inches
---
diagonal:
---
d = sqrt( aa + aa )
d = sqrt( 2aa )
d = sqrt( 2aa )
d = sqrt( 2 ) * sqrt( aa )
d = sqrt( 2 ) * 9
d = 1.41421356237 * 9
d = 12.7279220613 inches
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php